Helllllp I don't know how to find a discriminant. u_u
 ex: 2x^2+5x+6=0
 
 
 d=
 # of real number solutions:
 
 Also, suppose a parabola has a vertex in Quadrant I and a<0 in the equation y=ax^2+bx+c. How many real solutions will the equation ax^2+bx+c=o have?
 
 Answer by lawdej(58)
       
 here is the way i hope you will be satisfied
 for a quadratic equation ax^2+bx+c=0 the discriminant is b^2-4ac
 so here a=2,b=5,c=6
 thus b^2-4ac=5^2-(4x2x6)=25-48=-23 this indicates the equation has no solution
 in general if b^2-4ac<0 no solution,if b^2-4ac=0 one solution,if b^2-4ac>0 two solution
 | 
  
 | 
 |